OpenStudy (anonymous):

A cylindrical with a volume of 576πm^3 is circumscribed about a sq. prism which has one side of the base that measures 8m. what is the altitude of the cylinder?

OpenStudy (jamesj):

Draw a diagram of the cross-section and you'll see the radius of the circle must be \[ r = 4 \sqrt{2} \]

OpenStudy (anonymous):

i can't imagine the illustration

OpenStudy (jamesj):

 The length of each side of the square is 8; the length of the radius then is the hypothenuse of a right-triangle, where the other two sides are 4 and 4.

OpenStudy (anonymous):

i rearrange the formula and i got this - h = V ÷ (pi)r2.

OpenStudy (jamesj):

Yes, of course, as \[ V = \pi r^2 h \]

OpenStudy (jamesj):

V of the cylinder that is.

OpenStudy (anonymous):

so , my formula is right ?

OpenStudy (anonymous):

but i don't know how to get the radius of the circle.

OpenStudy (jamesj):

Yes, your formula is correct. For radius r, look again carefully at the diagram I've drawn above. r is the hypothenuse of a right-angled triangle, the other two sides being half of a side of the square.

OpenStudy (anonymous):

do i need to find the LSA of the cylinder? LSA=2pier^2

OpenStudy (anonymous):

h= v/πr^2 h= 576/π(4√2)^2 = 5.73m ?? is that right?

OpenStudy (jamesj):

No, V = 576.pi

OpenStudy (anonymous):

oh. i forgot . so h= 576π/π(4√2)^2 = 18m??

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

but i really don't get the formula for the radius